Required information [The following information applies to the questions displayed below.] The following transactions pertaina. Record the above transactions in a horizontal statements model like the following one. (in the Cash Flow column, indicateACCOUNTING SOLUTIONS INC. Horizontal Statements Model - 2020 Balance Sheet Income Statement Assets - Stockholders Equity Netb-1. Prepare an income statement for 2018, 2019, and 2020. ACCOUNTING SOLUTIONS INC. Income Statement For the Year Ended DeceACCOUNTING SOLUTIONS INC. Balance Sheet As of December 31 2018 2019 2020 Assets Total Assets Liabilities Stockholders equity
Answer #1

a.

Accounting Solutions Inc.
Horizontal Statements Model: 2018
Event 2018 Assets = Stockholders Equity Income Statement Statement of Cash Flows
Cash Equipment Accumulated Depreciation Common Stock Retained Earnings Revenue - Expense = Net Income
1 60,000 60,000 60,000 FA
2 (24,000) 24,000 (24,000) IA
3 (1,300) 1,300 (1,300) IA
4 39,600 39,600 39,600 39,600 OA
5 (1,060) (1,060) 1,060 (1,060 ) OA
6 10,120 (10,120) 10,120 - -
Bal. 73,240 25,300 (10,120) 60,000 28,420 39,600 11,180 28,420
Horizontal Statements Model:2019
2019
Bal. 73,240 25,300 (10,120) 60,000 28,420
1. (990) (990) 990 (990) OA
2. (1,080) (1,080) 1.080 (1,080) OA
3. 42,600 42,600 42,600 42,600 OA
4. (820) (820) 820 (820) OA
5. 6,072 (6,072) 6,072 - -
Bal. 112,950 25,300 (16,192) 60,000 62,058 42,600 8,962 33,638
Horizontal Statements Model : 2020
2020
Bal. 112,950 25,300 (16,192) 60,000 62,058
1 (3,100) 3,100 (3,100) IA
2 (910) (910) 910 (910) OA
3 31,590 31,590 31,590 31,590 OA
4. 6,104 (6,104) 6,104 - -
Bal. 140,530 28,400 (22,296) 60,000 86,634 31,590 7,014 24,576

Straight line depreciation rate = 100 / 5 = 20 %

Depreciation expense for 2018 = $ 25,300 x 20 % x 2 = $ 10,120

Depreciation expense for 2019 = $ (25,300 - 10,120) * 40 % = $ 6,072.

Depreciation expense for 2020 = [$ ( 25,300 - 16,192) + $ 3,100] x 100 / 4 * 2 = $ 6,104.